Exercise 16-36 (Static) Variable Cost Variances (LO 16-5)

Paynesville Corporation manufactures and sells a preservative used in food and drug manufacturing. The company carries no inventories. The master budget calls for the company to manufacture and sell 100,000 liters at a budgeted price of $75 per liter this year. The standard direct cost sheet for one liter of the preservative follows.

Direct materials (2 pounds @ $4) $ 8
Direct labor (0.5 hours @ $24) 12


Variable overhead is applied based on direct labor hours. The variable overhead rate is $20 per direct-labor hour. The fixed overhead rate (at the master budget level of activity) is $10 per unit. All non-manufacturing costs are fixed and are budgeted at $1.2 million for the coming year.
At the end of the year, the costs analyst reported that the sales activity variance for the year was $270,000 unfavorable.

The following is the actual income statement (in thousands of dollars) for the year.

Sales revenue $ 7,238
Less variable costs
Direct materials 748
Direct labor 1,010
Variable overhead 930
Total variable costs $ 2,688
Contribution margin $ 4,550
Less fixed costs
Fixed manufacturing overhead 1,050
Non-manufacturing costs 1,230
Total fixed costs $ 2,280
Operating profit $ 2,270

  During the year, the company purchased 176,000 pounds of material and employed 40,400 hours of direct labor.

Required;

a) Compute the direct material price and efficiency variances

b) Compute the direct labor price and efficiency variances

c) Compute the variable overhead price and efficiency variances
Answer #1

Solution:

Budgeted contribution margin per unit = Selling price - Variable cost per unit

= $75 - ($8 + $12 + $10) = $45 per unit

Actual units sold = Budgeted sales units - Unfavorable sales activity variance / Budgeted contribution margin per unit =  100000 - ($270,000 / $45) = 94000 liters

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
176000 $4.25 $7,48,000.00 176000 $4.00 $7,04,000.00 188000 $4.00 $7,52,000.00
$44,000.00 Unfavorable $48,000.00 Favorable
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $44,000.00 Unfavorable
Direct material quantity variance $48,000.00 Favorable
Total direct material variance $4,000.00 Favorable
Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
40400 $25.00 $10,10,000.00 40400 $24.00 $9,69,600.00 47000 $24.00 $11,28,000.00
$40,400.00 Unfavorable $1,58,400.00 Favorable
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $40,400.00 Unfavorable
Direct Labor Efficiency variance $1,58,400.00 Favorable
Total direct labor variance $1,18,000.00 Favorable
Variable Overhead Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
40400 $23.02 $9,30,000.00 40400 $20.00 $8,08,000.00 47000 $20.00 $9,40,000.00
$1,22,000.00 Unfavorable $1,32,000.00 Favorable
Variable overhead spending variance Variable overhead efficiency variance
Variable overhead Spending variance $1,22,000.00 Unfavorable
Variable overhead efficiency variance $1,32,000.00 Favorable
Variable overhad cost variance $10,000.00 Favorable